# Hidden Population III Descendants in Ultra-Faint Dwarf Galaxies

### v1

@@ 8 Conclusions
To investigate the chemical signatures left by Pop III SNe, we implement a theoretical chemical evolution model of Boötes I (Rossi et al., 2021, 2023), the best studied UFD galaxy. Our model accounts for the incomplete sampling of the IMF in this poorly star-forming system. The chemical enrichment from carbon to zinc is followed by including SNe and AGB stars from both Pop III stars and the subsequent Pop II stellar populations. Our key results are:
@@
  • Among literature data of Boötes I stars with $>5$ abundance ratios measured, we uncover one mono-enriched and two multi-enriched Pop III descendants (Fig.6). Specifically, our findings show that the CEMP star Boo-119 exhibits an abundance pattern consistent with enrichment by two faint Pop III SNe; Boo-94 formed in an environment polluted by three Pop III SNe (two ccSN and one high-energy SN); and, notably, we identify a mono-enriched Pop III hypernova descendant, Boo-130;
@@ 9 Acknowledgments
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 804240). S.S. and I.V. acknowledge support from the PRIN-MIUR2017, prot. n. 2017T4ARJ5. Á.S. acknowledges ERC support (grant agreement No. 101117455).
@@ Appendix A The probability to find Pop III descendants
The [C/Fe] ratio is commonly used as a diagnostic tool for identifying metal-poor stars enriched by Pop III sources. Additionally, the A(C) value is employed to discern the origins of distinct CEMP stellar populations: in Rossi et al. (2023) we show that among CEMP-no stars, true Pop III stars descendants appear for $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$.
To unveil if there exists a region in which Pop III descendants can be found, we derived the probability that a given stellar population has been imprinted by Pop III or Pop II SNe for different [Fe/H] and [C/Fe] (or A(C)) values. By inspecting the A(C)-[Fe/H] diagram in Fig.10 it is evident that two separate branches exist, both showing roughly constant A(C) values in the [Fe/H] range. The first branch, at $\rm A(C)\approx 7$ is dominated by stars predominantly enriched by Pop II stars. In the second branch, $\rm A(C)\lesssim 6$, the enrichment is driven by Pop III SNe. The two different branches are also evident in [C/Fe]-[Fe/H] space where we can note that Pop III and Pop II enriched stars are located in two parallel branches that exhibit a strong correlation with [Fe/H]. As already discussed in Rossi et al. 2023, Pop II enriched stars can be divided into two different groups with different chemical enrichment sources: stellar populations with $\rm[C/Fe]>+0.7$ and $\rm[Fe/H]\lesssim-4$, located in the branch at $\rm A(C)\gtrsim 7,$ are the so called moderate CEMP stars, which formed from an ISM mainly enriched by Pop II AGB stars. C-normal stars with $\rm[C/Fe]\lesssim+0.7$ and $\rm-4\lesssim[Fe/H]\lesssim-1$ form in birth clouds predominantly enriched by Pop II SNe. Indeed, Pop II SNe stellar yields adopted (Limongi & Chieffi, 2018), do not allow to form stellar populations enhanced in carbon, i.e. with $\rm[C/Fe]>+0.7$. Finally, carbon-enhanced stars with $\rm-3.5\lesssim[Fe/H]\lesssim-1$ formed out of gas enriched by both AGB and SNe Pop II stars.
Pop III enriched stars are located in the low-branch at $\rm A(C)\lesssim 6$ of the A(C)-[Fe/H] or equivalently in the bottom branch of the [C/Fe]-[Fe/H] diagram. The majority of Pop III stars descendants with $\rm[Fe/H]<-4$ are CEMP-no stars and the probability to find them increases as [Fe/H] decreases and it is maximum for $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$. Furthermore, it is predicted that Pop III descendants can be found among C-normal stars. In particular at $\rm[C/Fe]<0$ and $\rm-4\lesssim[Fe/H]\lesssim-1$ the probability to find the chemical signatures of pristine SNe exceeds $\gtrsim 50\%$. In conclusion, by exploiting both [C/Fe] and A(C) as diagnostics we can successfully predict promising descendants among both CEMP and C-normal stars: carbon-enhanced Pop III stars descendants are predicted to have $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$ while C-normal stars imprinted by Pop III SNe have $\rm[C/Fe]<0$ and $\rm[Fe/H]\gtrsim-3.5$. In other words we predict that stars with these properties have the highest probability of being Pop III descendants.

### v2

@@ 9 Conclusions
To investigate the chemical signatures left by Pop III SNe, we implement a theoretical chemical evolution model of Boötes I (Rossi et al., 2021, 2023), the best studied UFD galaxy (see also Romano et al. (2014); Vincenzo et al. (2014); Alexander et al. (2023)). Our model accounts for the incomplete sampling of the IMF in this poorly star-forming system. The chemical enrichment from carbon to zinc is followed by including SNe, Type Ia SNe (SN Ia) and AGB stars from both Pop III stars and the subsequent Pop II stellar populations. Our key results are:
@@
  • Among literature data of Boötes I stars with $>5$ abundance ratios measured, we uncover one mono-enriched and two multi-enriched Pop III descendants (Fig.6). Specifically, our findings show that the CEMP star Boo-119 exhibits an abundance pattern consistent with enrichment by two faint Pop III SNe; Boo-94 formed in an environment polluted by three Pop III SNe (one faint, one ccSN and one high-energy SN); and, notably, we identify a mono-enriched Pop III hypernova descendant, Boo-130;
@@ 10 Acknowledgments
This project has received funding from the European Research Council Executive Agency (ERCEA) under the European Union’s Horizon Europe research and innovation program (acronym TREASURES, grant agreement No 101117455”). M.R., S.S., I.V., and I.K. acknowledge ERC support (grant agreement No. 804240). S.S. and I.V. also acknowledge support from the PRIN-MIUR2017, prot. n. 2017T4ARJ5. We thank Stefano Ciabattini for insightful discussions and for sharing his new findings.
@@ Appendix A The probability to find Pop III descendants
The [C/Fe] ratio is commonly used as a diagnostic tool for identifying metal-poor stars enriched by Pop III sources . Additionally, the A(C) value is employed to discern the origins of distinct CEMP stellar populations (Bonifacio et al., 2015; Yoon et al., 2019): in Rossi et al. (2023) we show that among CEMP-no stars, true Pop III stars descendants appear for $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$.
To unveil if there exists a region in which Pop III descendants can be found, we derived the probability that a given stellar population has been imprinted by Pop III or Pop II SNe for different [Fe/H] and [C/Fe] (or A(C)) values. Analyzing the A(C)-[Fe/H] diagram in Fig. 10, a clear separation emerges between stars imprinted by Pop III and Pop II enrichment. Stars primarily enriched by Pop III SNe are characterized by $4\lesssim\rm A(C)\lesssim 8$ and are predominantly located at $\rm[Fe/H]<-3.5$. In contrast, stars primarily enriched by Pop II SNe exhibit $\rm[Fe/H]>-3.5$ and none of them fall within the CEMP star category.Indeed, Pop II SNe stellar yields adopted (Limongi & Chieffi, 2018), do not allow to form stellar populations enhanced in carbon, i.e. with $\rm[C/Fe]>+0.7$. Pop III enriched stars are located in the low-branch at $\rm A(C)\lesssim 6$ of the A(C)-[Fe/H] or equivalently in the bottom branch of the [C/Fe]-[Fe/H] diagram. The majority of Pop III stars descendants with $\rm[Fe/H]<-4$ are CEMP-no stars and the probability to find them increases as [Fe/H] decreases and it is maximum for $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$. Furthermore, it is predicted that Pop III descendants can be found among C-normal stars. In particular at $\rm[C/Fe]<0$ and $\rm-4\lesssim[Fe/H]\lesssim-1$ the probability to find the chemical signatures of pristine SNe exceeds $\gtrsim 50\%$. In conclusion, by exploiting both [C/Fe] and A(C) as diagnostics we can successfully predict promising descendants among both CEMP and C-normal stars: carbon-enhanced Pop III stars descendants are predicted to have $\rm A(C)\lesssim 6$ and $\rm[Fe/H]\lesssim-3.5$ while C-normal stars imprinted by Pop III SNe have $\rm[C/Fe]<0$ and $\rm[Fe/H]\gtrsim-3.5$. In other words we predict that stars with these properties have the highest probability of being Pop III descendants.
@@ Appendix B Pop II descendants
Appendix C Testing alternative Pop II stellar yields
In Fig.12 we compare the observed chemical abundance patterns of our three Pop III candidate stars (Boo-119, Boo-130, and Boo-94) with synthetic abundance patterns predicted from PopII-only enrichment scenarios. The models are computed using the stellar yield tables from Nomoto et al. (2013) at Z = -3, considering both standard core-collapse SNe and hypernovae. For each star, we test enrichment from either a single Pop II progenitor or a combination of two with different initial masses. The resulting chi-squared values are shown for each scenario and compared to those obtained using the PopIII-only enrichment in our fiducial model. In all cases, the PopIII models provide significantly better fits, supporting their identification as genuine PopIII descendants.
Figure 12: Comparison between observed abundance patterns of the three Pop III candidate stars Boo-119, Boo-130, and Boo-94 (colored points) and synthetic abundance patterns predicted from Pop II-only adopting the stellar yields from Nomoto et al. (2013) at $Z=-3$, including both standard core-collapse SN (NK, solid gray line) and hypernovae (HN, dashed gray line). The panels show the model that yields the lowest $\chi^{2}$ value in each case, reported in the legend along with the corresponding progenitor mass or mass pair. In all three cases, Pop II-only enrichment produces significantly worse fits compared to the Pop III-only fiducial model.